Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form Y=MX+b.
(-3,3);4x=5y+6
The equation of the line is Y=
(Simplify your answer. Use integers or fractions for any numbers in the expression)

Answer by mananth(16946)

(-3,3);4x=5y+6
-5y=-4x+6
Divide by-5
y=0.8x-1.2
Compare this equation with y=mx+b
slope m =0.8
The slope of a line parallel to the above line will be the same
The slope of the required line will be 0.8
m=0.8,point(-3, 3)
Find b by plugging the values of m & the point in
y=mx+b
3=-2.40 +b
b=5.4
m=0.8
Plug value of the slope and b
The required equation is y=0.8x+ 5.4